graph the line 2x-5y=-10
Answer by Fombitz(32388)

Quickest way is to find the intercepts and plot them.
To find the x-intercept, set y=0 and solve for x.
2x-5y=-10
2x=-10
x=-5
(-5,0)
To find the y-intercept, set x=0 and solve for y.
2x-5y=-10
-5y=-10
y=2
(0,2)
.
.
Plot those two points.
Then connect them with a line.
.
.